OpenStudy (abbycross167):

Can someone please check my work on this math question? please? I'll fan and give a medal!!

OpenStudy (abbycross167):

Use this equation to answer the question -4.9t^2 + 24.5t + 117.6 = 0 determine when a ball would be at the height of 49 meters

jimthompson5910 (jim_thompson5910):

The height function is h(t) = -4.9t^2 + 24.5t + 117.6 replace h(t) with 49 to get 49 = -4.9t^2 + 24.5t + 117.6

jimthompson5910 (jim_thompson5910):

so your goal is to solve 49 = -4.9t^2 + 24.5t + 117.6 for t get everything to one side and then use the quadratic formula

OpenStudy (abbycross167):

so would I put 49 in the equation or just leave it out? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

Subtract 49 from both sides

OpenStudy (abbycross167):

oh ok

jimthompson5910 (jim_thompson5910):

49 = -4.9t^2 + 24.5t + 117.6 49-49 = -4.9t^2 + 24.5t + 117.6-49 0 = -4.9t^2 + 24.5t + 68.6 Then use the quadratic formula

OpenStudy (abbycross167):

x = - (-24.5) +/- \/([-24.5]2 - 4*4.9*-68.6) 2(4.9) x = 24.5 +/- \/1944.81 9.8 T=6.999999999999999, -2

jimthompson5910 (jim_thompson5910):

I'm not sure why your calculator isn't saying 7 and instead it says 6.999999999999999

jimthompson5910 (jim_thompson5910):

Anyways, you should get t = 7 or t = -2. We ignore t = -2 if t is a time value (which I'm assuming it is) because negative time values make no sense. So the object reaches 49 meters at t = 7 seconds

OpenStudy (abbycross167):

thank you so much!!!

jimthompson5910 (jim_thompson5910):

you're welcome